The rate constant for a reaction of zero order in A is 0.0030 mol L-1 s-1. How long will it take for the initial concentration of A to fall from 0.10 M to 0.075 M?

Grade:

1 Answers

Sunil Kumar FP
9 years ago
The rate constant for a reaction of zero order in A is 0.0030 mol L-1 s-1. Time will it take for the initial concentration of A to fall from 0.10 M to 0.075 M
The rate constant for 0 order reaction is

[A]=[A]0−kt

here,A=.075 M and A0 =.1M
A0 -A= .1-.075 =.025M
time taken =A0 -A/k =.025/.0030 s =8.3 second